OpenStudy (anonymous):

Simplify 1/1+sin theta + 1/1-sin theta

OpenStudy (akashdeepdeb):

\[\frac{1}{1+sin~\theta} + \frac{1}{1-sin~\theta}\] Multiply and divide both terms by their conjugate [To make the denominators same] \[\frac{1 * (1-sin\theta)}{(1+sin~\theta)(1-sin\theta)} + \frac{1 * (1+sin\theta)}{(1-sin~\theta)(1+sin\theta)}\] Now you've got the denominators to be equal! Then use these rules: \(1-sin^2\theta = cos^2\theta ~~~AND~~~\frac{1}{cos\theta} = sec\theta\) After simplifying more you'll get: \[\frac{1-sin\theta}{1-sin^2~\theta} + \frac{1+sin\theta}{1-sin^2~\theta} = \frac{2}{cos^2\theta} = 2sec^2\theta\]